A company issued 10.0%, 5-year bonds with a par value of $240,000. The market rate when the bonds were issued was 11.0%. The company received $230,954.85 cash for the bonds. Using the effective interest method, the amount of interest expense for the second semiannual interest period is:

Multiple Choice

  • $12,000.00.

  • $12,741.16.

  • $12,702.52.

  • $24,000.00.

  • $25,443.68.
Answer #1
Interest expense for first semiannual interest period 12702.52 =230954.85*11%/2
Cash interest paid 12000 =240000*10%/2
Discount amortized for first semiannual interest period 702.52 =12702.52-12000
Carrying value after first semiannual interest period 231657.37 =230954.85+702.52
Interest expense for the second semiannual interest period 12741.16 =231657.37*11%/2
Option B $12,741.16 is correct